# suppSupplementary References

### v1

@@ 3.2 Performance and Memory
Throughput. Table 2 shows the throughput (i.e., signals processed per second per CPU thread) that Rawsamble reports. Our goal is to estimate the number of CPU threads needed to achieve a throughput faster than the throughput of a single sequencer. Using as few CPU threads as possible is useful to 1) provide better scalability (i.e., analyzing a larger amount of data with the same computation capabilities) and 2) reduce the overall computational requirements and corresponding energy consumption (i.e., analyzing the same amount of data with less computational capabilities). The latter is especially essential for resource-constrained devices (e.g., devices with external and limited batteries). The throughput of a single nanopore is around 5,000 signals per second [35], and the entire sequencer is usually equipped with 512 nanopores. This means a single sequencer’s throughput is around $2{,}560{,}000$ signals per second ($5{,}000\times 512$[85]. We find that Rawsamble provides an average throughput of $2{,}432{,}561$ signals per second per CPU thread. This means Rawsamble can achieve an analysis throughput faster than a single sequencer’s throughput by using an average of two CPU threads. This shows that the overlapping mechanism of Rawsamble is fast enough for performing real-time overlapping tasks using very few CPU threads.
@@
D1 D2 D3 D4 D5
SARS-CoV-2 E. coli Yeast Green Algae Human
Throughput 2,065,764 2,720,702 2,128,800 1,668,065 3,579,472
Computational resources. Table 3 shows the computational resources of the different approaches. Inside the parentheses provided with the minimap2 results, we provide the ratio between the reported result and the corresponding Rawsamble result (if higher than 1×1$\times$1 ×, Rawsamble is better).
Table 3: Comparison of various tools across different organisms in terms of elapsed time, CPU time, and peak memory usage. The values in parentheses represent the ratio of the result shown in the cell compared to the corresponding result of Rawsamble (values higher than 1×1$\times$1 × indicate that Rawsamble performs better). We highlight the cells that Rawsamble provide a better and worse results with colors.
@@
D1 Rawsamble 0:00:03 33 1.07
SARS-CoV-2 Minimap2 0:00:01 (0.33×0.$3$) 19 (0.58×0.58$\times$0.58 ×) 0.16 (0.15×0.15$\times$0.15 ×)
Minimap2 + Dorado CPU (Fast) 0:01:45 (35.00×35.$0$) 3,227 (97.79×97.79$\times$97.79 ×) 44.93 (41.99×41.$9$)
Minimap2 + Dorado CPU (HAC) 0:05:45 (115.00×115.$0$) 5,457 (165.36×165.36$\times$165.36 ×) 57.98 (54.19×54.19$\times$54.19 ×)
Minimap2 + Dorado GPU (HAC) 0:01:41 (33.67×33.67$\times$33.67 ×) NA 0.8 (0.75×0.75$\times$0.75 ×)
Minimap2 + Dorado GPU (SUP) 0:25:47 (515.67×515.67$\times$515.67 ×) NA 1.23 (1.15×1.15$\times$1.15 ×)
D2 Rawsamble 1:12:44 132,758 6.72
E. coli Minimap2 0:14:25 (0.20×0.20$\times$0.20 ×) 25,721 (0.19×0.19$\times$0.19 ×) 26.73 (3.98×3.98$\times$3.98 ×)
Minimap2 + Dorado CPU (Fast) 7:17:05 (6.01×6.01$\times$6.01 ×) 583,358 (4.39×4.39$\times$4.39 ×) 50.43 (7.50×7.50$\times$7.50 ×)
Minimap2 + Dorado CPU (HAC) 32:26:12 (26.76×26.76$\times$26.76 ×) 1,335,697 (10.06×10.06$\times$10.06 ×) 38.0 (5.65×5.65$\times$5.65 ×)
Minimap2 + Dorado GPU (HAC) 0:36:14 (0.50×0.50$\times$0.50 ×) NA 26.73 (3.98×3.98$\times$3.98 ×)
Minimap2 + Dorado GPU (SUP) 1:30:30 (1.24×1.24$\times$1.24 ×) NA 26.73 (3.98×3.98$\times$3.98 ×)
D3 Rawsamble 0:01:18 2,241 6.39
Yeast Minimap2 0:00:21 (0.27×0.27$\times$0.27 ×) 290 (0.13×0.13$\times$0.13 ×) 5.25 (0.82×0.82$\times$0.82 ×)
Minimap2 + Dorado CPU (Fast) 0:54:04 (41.59×41.59$\times$41.59 ×) 71,796 (32.04×32.04$\times$32.04 ×) 56.13 (8.78×8.78$\times$8.78 ×)
Minimap2 + Dorado CPU (HAC) 3:13:56 (149.18×149.18$\times$149.18 ×) 193,640 (86.41×86.41$\times$86.41 ×) 65.43 (10.24×10.24$\times$10.24 ×)
Minimap2 + Dorado GPU (HAC) 0:04:33 (3.50×3.50$\times$3.50 ×) NA 5.25 (0.82×0.82$\times$0.82 ×)
Minimap2 + Dorado GPU (SUP) 0:10:33 (8.12×8.12$\times$8.12 ×) NA 5.92 (0.93×0.93$\times$0.93 ×)
D4 Rawsamble 0:07:57 14,064 8.67
Green Algae Minimap2 0:00:47 (0.10×0.10$\times$0.10 ×) 882 (0.06×0.06$\times$0.06 ×) 8.7 (1.00×1.$0$)
Minimap2 + Dorado CPU (Fast) 1:16:35 (9.63×9.63$\times$9.63 ×) 79,606 (5.66×5.$6$) 50.88 (5.87×5.87$\times$5.87 ×)
Minimap2 + Dorado CPU (HAC) 4:30:07 (33.98×33.98$\times$33.98 ×) 286,362 (20.36×20.36$\times$20.36 ×) 64.07 (7.39×7.39$\times$7.39 ×)
Minimap2 + Dorado GPU (HAC) 0:06:01 (0.76×0.76$\times$0.76 ×) NA 8.7 (1.00×1.$0$)
Minimap2 + Dorado GPU (SUP) 0:14:54 (1.87×1.87$\times$1.87 ×) NA 8.7 (1.00×1.$0$)
D5 Rawsamble 0:28:56 51,975 6.0
Human Minimap2 0:01:52 (0.06×0.06$\times$0.06 ×) 1,372 (0.03×0.03$\times$0.03 ×) 20.21 (3.37×3.37$\times$3.37 ×)
Minimap2 + Dorado CPU (Fast) 6:42:24 (13.91×13.91$\times$13.91 ×) 802,983 (15.45×15.45$\times$15.45 ×) 81.98 (13.66×13.$6$)
Minimap2 + Dorado CPU (HAC) 23:27:18 (48.64×48.64$\times$48.64 ×) 1,219,043 (23.45×23.45$\times$23.45 ×) 46.12 (7.69×7.69$\times$7.69 ×)
Minimap2 + Dorado GPU (HAC) 0:20:24 (0.71×0.71$\times$0.71 ×) NA 20.31 (3.38×3.38$\times$3.38 ×)
Minimap2 + Dorado GPU (SUP) 1:05:48 (2.27×2.27$\times$2.27 ×) NA 20.21 (3.37×3.37$\times$3.37 ×)
@@
We make three key observations. First, Rawsamble provides a substantial speedup and lower peak memory usage compared to minimap2 when combined with Dorado’s fast model on a CPU by, on average, 16.36×16.36$\times$16.36 × (elapsed time), 16.45×16.45$\times$16.45 × (CPU time), and 11.73×11.73$\times$11.73 × (peak memory usage). When using Dorado’s HAC model, Rawsamble provides even better results on average by 59.70×59.70$\times$59.70 × (elapsed time), 36.93×36.93$\times$36.93 × (CPU time), and 11.23×11.23$\times$11.23 × (peak memory usage), since running the HAC model is computationally more costly than running the fast model.
@@
Second, Rawsamble achieves an average: 1) speedup of 1.99×1.$9$ (HAC) and 7.40×7.40$\times$7.40 × (SUP), and 2) reduced peak memory usage of 1.53×1.53$\times$1.53 × (HAC) and 1.70×1.70$\times$1.70 × (SUP), compared to GPU-based basecalling followed by minimap2. Although basecalling with GPUs followed by overlapping is usually faster in most cases when using the HAC model (Minimap2 + Dorado GPU (HAC) in Table 3), Rawsamble still achieves better average speedup due to the substantial speedup it provides for the D1 dataset. Although comparing the results between CPUs and GPUs is not ideal due to the massive parallelism that GPUs provide compared to CPUs, Rawsamble still provides comparable and even better performance even with the limited parallelism (i.e., 32 threads) it uses on CPUs compared to the parallelism that GPUs provide (e.g., a few tens of thousands of threads). This shows that Rawsamble can provide even faster results when accelerated with GPUs (outside the scope of this work) based on its comparison when basecalling is done using CPUs and GPUs.
Third, minimap2, without considering the resources that basecalling requires, is more resource-efficient than Rawsamble as it takes on average 0.16×0.16$\times$0.16 × (elapsed time), 0.12×0.12$\times$0.12 × (CPU time), and $1\times1$ (peak memory usage) of those Rawsamble takes. This is mainly expected as analyzing raw signals requires additional steps (e.g., signal-to-event conversion), more memory due to sequencing depth (as opposed to a reference genome), and incurs noise that requires less efficient parameter settings for Rawsamble (e.g., the window parameter in minimizers) when analyzing raw signals compared to minimap2. Future work can focus on processing these raw signals more accurately to reduce the limits for adjusting parameters that can impact performance and memory usage.
@@ 3.3 All-vs-All Overlapping Statistics
Table 4 shows the all-vs-all overlapping statistics between Rawsamble and minimap2, where each row sums to 100%. We make two key observations. First, on average, $36.57\%$ of overlap pairs generated by Rawsamble are shared with the overlap pairs that minimap2 generates. This shows that a large fraction of overlapping pairs does not require basecalling to generate identical overlapping information identified by minimap2 after basecalling. Second, although Rawsamble can find a substantial amount of overlap pairs identical to the overlaps minimap2 finds, specifically for the D1 dataset, there are still overlapping pairs unique to Rawsamble ($16.25\%$ on average) and minimap2 ($47.17\%$ on average), mainly due to the decreased shared overlaps as the genome size increases. These differences are likely due to differences in 1) certain parameters (e.g., chaining scores) and 2) increased noise inherent in raw signals compared to basecalled sequences, which can become more pronounced in genomes with greater size, complexity, or repetitiveness. Additionally, as observed in an earlier work [86], constructing contiguous assemblies is not dependent on finding as many overlapping pairs as possible. Instead, contiguous assemblies can still be constructed using a smaller but useful portion of overlapping pairs. We conclude that Rawsamble provides a mechanism that shares a large portion of overlaps with minimap2, while the shared portions decrease as the genome size increases.
@@
D1 SARS-CoV-2 11.55 15.27 73.18
D2 E. coli 8.33 50.62 41.05
D3 Yeast 24.94 35.17 39.89
D4 Green Algae 3.76 78.64 17.61
D5 Human 32.69 56.18 11.13
3.4 De novo Assembly Evaluations
To evaluate the impact of differing reported overlaps between Rawsamble and minimap2, we construct de novo assemblies from these overlaps. Table 5 shows the statistics of these de novo assemblies. We do not provide the assembly statistics for the D1 dataset as miniasm cannot generate an assembly using the overlaps from Rawsamble and minimap2, potentially due to the small size of the genome. We make three observations.
@@ 3.4 De novo Assembly Evaluations
First, we find that the overlaps found using Rawsamble can form long paths during assembly that lead to assembly constructions and large connected components. Rawsamble is the first work that enables de novo assembly construction directly from raw signal overlaps without basecalling, which has several implications and can enable future work, as we discuss in Section 4.
@@
Second, we observe that the unitigs generated from Rawsamble overlaps are usually less contiguous compared to those generated from minimap2 overlaps, based on all the metrics we show in Table 5 and Supplementary Figures S2S5. Compared to the gold standard assemblies generated using highly accurate basecalled reads and a state-of-the-art assembler, we find that Rawsamble can still achieve a significant portion of the assembly contiguity, especially for the D2 dataset (E.coli). For example, Rawsamble constructs unitigs with the longest unitig length of 2.7 million bases, which is over half the length of the E. coli genome, whereas the gold standard assembly produces a single unitig covering the entire genome. This indicates that Rawsamble can achieve substantial contiguity relative to the gold standard without basecalling.
Third, for the larger genomes in the D4 and D5 datasets, we find that Rawsamble can construct longer unitigs compared to those generated by minimap2 and, notably, achieves better contiguity than the gold standard assemblies in the D5 dataset. The coverage of these datasets is substantially lower than that of other datasets. This suggests that Rawsamble can find more useful overlaps leading to longer paths in the assembly for the cases when coverage is relatively lower (e.g., coverage of less than or around 5×5$\times$5 × in Table 1). Basecalled sequences may generate more contiguous assemblies when the coverage is higher. We conclude that Rawsamble generates useful overlapping information, enabling the construction of assemblies directly from raw signals. This approach achieves substantial contiguity and, in some cases (e.g., at low coverage), provides superior contiguity compared to minimap2 and gold standard assemblies.
Table 5: Assembly Statistics.
Dataset Tool Total Largest N50 auN Longest Unitig
Length (bp) Comp. (bp) (bp) (bp) Unitig (bp) Count
D2 Rawsamble 14,525,505 4,841,669 1,535,079 1,309,738 2,722,499 31
E. coli minimap2 10,434,542 5,207,206 5,204,754 5,194,738 5,207,206 4
Gold standard 5,235,343 5,235,343 5,235,343 5,235,343 5,235,343 1
D3 Rawsamble 13,898,208 362,050 41,118 48,106 161,883 396
Yeast minimap2 23,755,455 1,611,876 134,050 150,908 464,054 282
Gold standard 11,963,521 11,835,059 640,934 623,210 1,073,346 68
D4 Rawsamble 3,448,899 448,422 93,111 108,818 252,038 50
Green Algae minimap2 2,117,190 198,709 63,310 88,906 198,709 55
Gold standard 106,479,288 2,255,807 452,774 538,136 1,667,975 420
D5 Rawsamble 1,850,419 493,004 51,300 116,049 364,113 48
Human minimap2 747,607 65,951 19,476 22,103 48,424 61
Gold standard 8,365,210 367,305 19,329 29,697 150,470 592

### v2

@@ 3.2 Performance and Memory
Throughput. Table 2 shows the throughput (i.e., signals processed per second per CPU thread) that Rawsamble reports. Our goal is to estimate the number of CPU threads needed to achieve a throughput faster than the throughput of a single sequencer. Using as few CPU threads as possible is useful to 1) provide better scalability (i.e., analyzing a larger amount of data with the same computation capabilities) and 2) reduce the overall computational requirements and corresponding energy consumption (i.e., analyzing the same amount of data with less computational capabilities). The latter is especially essential for resource-constrained devices (e.g., devices with external and limited batteries). The throughput of a single nanopore is around 5,000 signals per second [35], and the entire sequencer is usually equipped with 512 nanopores. This means a single sequencer’s throughput is around $2{,}560{,}000$ signals per second ($5{,}000\times 512$[81]. We find that Rawsamble provides an average throughput of $1{,}496{,}148$ signals per second per CPU thread. This means Rawsamble can achieve an analysis throughput faster than a single sequencer’s throughput by using an average of two CPU threads. This shows that the overlapping mechanism of Rawsamble is fast enough for performing real-time overlapping tasks using very few CPU threads.
@@
D1 D2 D3 D4
E. coli Yeast Green Algae E. coli (R10.4.1)
Throughput 1,894,434 1,304,910 1,668,065 1,117,183
Computational resources. Table 3 shows the computational resources of the different approaches. Inside the parentheses provided with Dorado + Minimap2 results, we provide the ratio between the reported result and the corresponding Rawsamble result (if higher than 1×1$\times$1 ×, Rawsamble is better).
Table 3: Comparison of various tools across different organisms in terms of elapsed time, CPU time, and peak memory usage. The values in parentheses represent the ratio of the result shown in the cell compared to the corresponding result of Rawsamble (values higher than 1×1$\times$1 × indicate that Rawsamble performs better). We highlight the cells that tools provide a better and worse results with colors.
@@
D1 Rawsamble 0:52:07 184,938 7.41
E. coli Dorado CPU (Fast) + Minimap2 2:56:45 (3.39×3.39$\times$3.39 ×) 593,764 (3.21×3.21$\times$3.21 ×) 36.73 (4.96×4.96$\times$4.96 ×)
Dorado CPU (HAC) + Minimap2 13:31:19 (15.57×15.57$\times$15.57 ×) 1,408,712 (7.62×7.62$\times$7.62 ×) 62.34 (8.41×8.41$\times$8.41 ×)
Dorado GPU (HAC) + Minimap2 0:26:48 (0.51×0.51$\times$0.51 ×) NA 26.73 (3.61×3.61$\times$3.61 ×)
D2 Rawsamble 0:01:21 3,753 6.68
Yeast Dorado CPU (Fast) + Minimap2 0:31:11 (23.10×23.10$\times$23.10 ×) 98,558 ($\times26$) 146.95 (22.00×22.$0$)
Dorado CPU (HAC) + Minimap2 2:30:28 (111.46×111.46$\times$111.46 ×) 548,495 (146.15×146.15$\times$146.15 ×) 290.85 (43.54×43.54$\times$43.54 ×)
Dorado GPU (HAC) + Minimap2 0:02:00 (1.48×1.48$\times$1.48 ×) NA 5.98 (0.90×0.90$\times$0.90 ×)
D3 Rawsamble 0:06:20 21,769 10.87
Green Algae Dorado CPU (Fast) + Minimap2 0:30:17 (4.78×4.78$\times$4.78 ×) 103,102 (4.74×4.74$\times$4.74 ×) 38.18 (3.51×3.51$\times$3.51 ×)
Dorado CPU (HAC) + Minimap2 7:29:20 (70.95×70.95$\times$70.95 ×) 178,106 (8.18×8.18$\times$8.18 ×) 20.76 (1.91×1.91$\times$1.91 ×)
Dorado GPU (HAC) + Minimap2 0:03:04 (0.48×0.48$\times$0.48 ×) NA 10.08 (0.93×0.93$\times$0.93 ×)
D4 Rawsamble 16:04:39 2,672,608 112.31
E. coli Dorado CPU (Fast) + Minimap2 17:16:52 (1.07×1.07$\times$1.07 ×) 3,478,136 (1.30×1.30$\times$1.30 ×) 103.1 (0.92×0.92$\times$0.92 ×)
(R10.4.1) Dorado CPU (HAC) + Minimap2 135:05:47 (8.40×8.40$\times$8.40 ×) 19,051,635 (7.13×7.13$\times$7.13 ×) 103.1 (0.92×0.92$\times$0.92 ×)
Dorado GPU (HAC) + Minimap2 4:59:49 (0.31×0.31$\times$0.31 ×) NA 103.1 (0.92×0.92$\times$0.92 ×)
@@
We make three key observations. First, Rawsamble provides a substantial speedup and lower peak memory usage compared to Dorado’s fast model on a CPU followed by minimap2 (Dorado CPU (Fast) + Minimap2) by, on average, 4.48×4.48$\times$4.48 × (elapsed time), 4.77×4.$7$ (CPU time), and 4.33×4.$3$ (peak memory usage). When using Dorado’s HAC model on a CPU, Rawsamble provides even better results on average by 31.89×31.89$\times$31.89 × (elapsed time), 15.96×15.96$\times$15.96 × (CPU time), and 5.03×5.03$\times$5.03 × (peak memory usage), since running the HAC model is computationally more costly than running the fast model.
@@
Second, GPU-based basecalling followed by minimap2 takes 0.58×0.58$\times$0.58 × of the time that Rawsamble takes while requiring higher peak memory usage by 1.29×1.29$\times$1.29 × compared to Rawsamble. Although comparing the results between CPUs and GPUs is not ideal due to the massive parallelism that GPUs provide compared to CPUs, Rawsamble still provides comparable performance given the limited parallelism it uses on CPUs (i.e., 64 threads) compared to the massive parallelism that GPUs provide (e.g., a few tens of thousands of threads). This shows that Rawsamble can provide even faster results when accelerated with GPUs based on its comparison when basecalling is done using CPUs and GPUs, which we leave as future work.
@@ 3.3 All-vs-All Overlapping Statistics
Table 4 shows the all-vs-all overlapping statistics between Rawsamble and minimap2. We make two key observations. First, on average, $30.94\%$ of the overlap pairs generated by Rawsamble are shared with the overlap pairs that minimap2 generates. This shows that a large fraction of overlapping pairs does not require basecalling to generate identical overlapping information identified by minimap2 after basecalling. Second, although Rawsamble can find a substantial amount of overlap pairs identical to the overlaps minimap2 finds, there are still overlapping pairs unique to Rawsamble ($9.67\%$ on average) and minimap2 ($59.39\%$ on average). These differences are likely due to differences in 1) certain parameters (e.g., chaining scores) and 2) increased noise inherent in raw signals compared to basecalled sequences, which can become more pronounced in genomes with greater size, complexity, or repetitiveness. Although maximizing the shared overlaps can provide insights regarding the accuracy of overlapping information that Rawsamble finds, it is not necessary to provide near-identical shared overlap statistics for certain use cases such as constructing de novo assemblies [82]. Instead, contiguous and more accurate assemblies can still be constructed using a smaller but useful portion of shared overlapping pairs [82]. We conclude that Rawsamble provides a mechanism that shares a large portion of overlaps with minimap2, while the shared portions decrease due to the differences in parameters and the noise in raw signal datasets.
@@
D1 E. coli 8.33 50.62 41.05
D2 Yeast 26.54 28.62 44.84
D3 Green Algae 3.76 78.64 17.61
D4 E. coli (R10.4.1) 0.06 79.69 20.25
3.4 Contiguity and Estimated Accuracy of the Assembly Graphs
To evaluate the impact of overlaps that Rawsamble and minimap2 find, we construct de novo assembly graphs from these overlaps by using miniasm off-the-shelf. Table 5 shows the contiguity statistics and estimated accuracy of these de novo assembly graphs (please see Section 3.1 for our explanation and discussion related to using these accuracy ratios only for comparison purposes, not for calculating the actual assembly identities). We note that the assembly graphs constructed using miniasm provide the connection information used to assemble unitigs, as well as the connections between unitigs that make up subgraph components in the assembly graph. Since we use the miniasm assembler without any modifications, we cannot generate the assembled sequences from raw signals as miniasm supports generating assembled sequences from basecalled sequencing reads. This mainly prevents us from evaluating the actual identity of these assemblies from raw signals, which we further discuss in Section 4.
Table 5: Assembly Statistics. Chained Read Percentage provides our estimate in terms of assembly graph accuracy. Other columns provide our evaluations in terms of assembly contiguity. For the accuracy, we highlight the cells that tools provide a better and worse results with colors.
Dataset Tool Chained Read Largest N50 auN Longest Unitig
Percentage (%) Comp. (bp) (bp) (bp) Unitig (bp) Count
D1 Rawsamble 34.0 4,841,669 1,535,079 1,187,229 2,347,310 32
E. coli minimap2 24.3 5,207,206 5,204,754 5,194,495 5,207,206 4
Gold standard NA 5,235,343 5,235,343 5,235,343 5,235,343 1
D2 Rawsamble 43.5 362,050 41,118 48,106 161,883 396
Yeast minimap2 43.2 1,611,876 134,050 137,172 464,054 282
Gold standard NA 11,835,059 640,934 623,210 1,073,346 68
D3 Rawsamble 41.3 448,422 93,111 107,914 252,038 50
Green Algae minimap2 53.2 198,709 63,310 91,360 198,709 55
Gold standard NA 2,255,807 452,774 538,136 1,667,975 420
D5 Rawsamble 34.6 22,921,373 27,460 43,505 410,905 2,194
E. coli minimap2 16.3 27,250,402 24,111 814,807 5,117,413 2888
(R10.4.1) Gold standard NA 5,230,610 5,230,610 5,230,610 5,230,610 1
@@ 3.4 Contiguity and Estimated Accuracy of the Assembly Graphs
First, we find that we can construct long unitigs from the raw signal overlaps that Rawsamble finds. The unitigs we can construct from these raw signal overlaps are substantially longer than the average read length of their corresponding datasets (e.g., the longest unitig length in D1 is 413.25×413.25$\times$413.25 × longer than the average read length of the D1 dataset). Rawsamble is the first work that enables de novo assembly construction directly from raw signal overlaps without basecalling, which has several implications and can enable future work, as we discuss in Section 4.
@@
Second, we observe that the unitigs generated from Rawsamble overlaps are usually less contiguous compared to those generated from minimap2 overlaps, based on all the metrics we show in Table 5. Compared to the gold standard assemblies generated using highly accurate basecalled reads and a state-of-the-art assembler, we find that Rawsamble can still achieve a significant portion of the assembly contiguity, especially for the D1 dataset (E.coli). For example, Rawsamble constructs unitigs with the longest unitig length of 2.3 million bases, which is almost half the length of the E. coli genome, whereas the gold standard assembly produces a single unitig covering the entire genome. This indicates that Rawsamble can achieve substantial contiguity relative to the gold standard without basecalling.
Third, in terms of the accuracy estimate of the assembly graphs (i.e., Chained Read Percentage), we find that a larger percentage of unitig reads generated by the Rawsamble overlaps appear in close proximity when these reads are mapped to their corresponding reference genome compared to the unitig reads generated by the minimap2 overlaps. Although this observation by itself does not show that the Rawsamble overlaps can lead to assemblies with higher identity, it provides the accuracy estimate that these assemblies share similar accuracy to those generated by the minimap2 overlaps.
We conclude that Rawsamble generates useful overlapping information, enabling the construction of assemblies directly from raw signals. The assembly graphs generated by the Rawsamble overlaps mainly provide comparable contiguity and accuracy compared to those generated by minimap2. We discuss the potential next steps to enable evaluating the accuracy of these assemblies in Section 4.